The difference between two numbers is 28. Three times the smaller is equal to 4 more than the larger. What are the numbers?
Answer by addingup(3677)

x-y = 28 add y on both sides:
x = 28+y
``````````````````````````````
3x = y+4 substitute for x:
3(28+y) = y+4
84+3y = y+4
80 = -2y
y = -40 this is one number
-40+28 = -12 this is the other number
-40+4 = 3(-12)
-36 = -36 I have the correct answer